OpenStudy (anonymous):

A local college student will be randomly selected to attend a leadership conference in Washington, D.C. There are 2060 local college students, of which 40% are male. 520 of the college students are graduate students, and half of the graduate students are female. Which of the following best represents the probability that the selected college student will be a male student or a graduate student?

OpenStudy (anonymous):

half of the graduate students are female, so presumably half are male how many male graduate students are the?

OpenStudy (anonymous):

250

OpenStudy (anonymous):

hmm what is half of 520?

OpenStudy (anonymous):

260 right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

now what is 40% of 2060?

OpenStudy (anonymous):

824

OpenStudy (anonymous):

ok so we have 824 male students all together, and another \(260\) female graduate students, for a total of \(824+260=1084\) students that are either male OR female graduate students

OpenStudy (anonymous):

your probability is therefore \[\frac{1084}{2060}\]

OpenStudy (anonymous):

Omg thank you so much!!

OpenStudy (anonymous):

which reduces to \[\frac{271}{515}\] yw